OpenStudy (camzzzie):

Multiple Choice 1. The sides of a triangle are in the ratio 3 : 4 : 5. What is the length of each side if the perimeter of the triangle is 90 cm? (1 point) (1 pt) 22.5 cm, 30 cm, and 37.5 cm (0 pts) 19.3 cm, 25.7 cm, and 32.1 cm (0 pts) 7.5 cm, 11.5 cm, and 32.1 cm (0 pts) 10.5 cm, 11.5 cm, and 12.5 cm 1 /1 point 2. Is the following always, sometimes, or never true? 6 + 8x – 9 = 11x + 14 – 3x (1 point) (0 pts) always true (0 pts) sometimes true (1 pt) never true 0 /1 point 3. Solve the inequality. Graph the solution set. 5r + 4 ≤ 5 (1 point) (1 pt) (0 pts) (0 pts) (0 pts) 4

OpenStudy (camzzzie):

4. Is the following inequality sometimes, always, or never true? 2(8x – 4) – 2x ≤ 14x + 12 (1 point) (0 pts) sometimes true (1 pt) always true (0 pts) never true 1 /1 point 5. Solve the compound inequality. Graph the solution. 6x – 4 < –34 or 3x + 10 > 10 (1 point) (0 pts) (1 pt) (0 pts) (0 pts) 0 /1 point 6. –2 ≤ 2x – 4 < 4 (1 point) (0 pts) (0 pts) (1 pt) (0 pts) 1 /1 point

OpenStudy (camzzzie):

7. Solve the absolute value equation. Graph the solution. 2|3x – 5| – 8 = 8 (1 point) (0 pts) (1 pt) (0 pts) (0 pts) 1 /1 point 8. Solve the equation. Check for extraneous solutions. 6|6 – 4x| = 8x + 4 (1 point) (0 pts) x = 1 (0 pts) x = or x = 2 (0 pts) x = (1 pt) x = or x = 1 0 /1 point 9. Solve the inequality. Graph the solution. (1 point) (0 pts)

OpenStudy (camzzzie):

(0 pts) (1 pt) (0 pts) 0 /1 point 10. A furniture maker uses the specification 19.88 ≤ w ≤ 20.12 for the width w in inches of a desk drawer. Write the specification as an absolute value inequality. (1 point) (0 pts) |w – 0.24| ≤ 20.12 (1 pt) |w – 20| ≤ 0.12 (0 pts) |w – 20| ≤ 0.24 (0 pts) |w – 0.12| ≤ 20 1 /1 point

OpenStudy (camzzzie):

this was hard for me so I hope this helps someone

OpenStudy (anonymous):

Do you take ALG 2A at connections academy?

OpenStudy (turtle_man):

Thank you so much!